# The tau function for ABS equations James Atkinson, October 28, 2024

### v1

@@ 1.1 Example
$u=np+mq,\quad p^{2}=\alpha,q^{2}=\beta.$ (5)
@@ 3 Equations for tau
which is equivalent to the form in [5]. Bilinear relations are given later, but they require the introduction of a non-autonomous scaling factor (gauge factor). The most direct statement of the more refined higher dimensional characterisation uses an intermediary Schwarzian variable [6].
@@
Like $\tau$, the quantity $\bar{\tau}/{\underaccent{\bar}{\tau}}$ is also not invariant under fractional-linear point transformations of $u$. An invariant Schwarzian variable [6] is
@@
where $w$ is any given singular solution of the quad equation. $\varphi$ in (30) is characterised by equations that are invariant under fractional linear transformations, and satisfies the lattice Schwarzian KP equation [8, 9] in $\mathbb{Z}^{3}$. If $w$ is constant, then $\varphi$ satisfies the lattice Schwarzian KdV equation, and all equations in Table 1 have a singular solution that is constant throughout $\mathbb{Z}^{d}$, with the exception of $A2$ and $Q4$. The canonical forms in Table 1 are such that the constant singular solution is $w=\infty$ in all cases, so substituting this into (30) offers an explanation for the isolation of $\tau$ leading to Proposition 3.2.
@@ 4 Reduced equations for tau
in term of the Hirota variable,
@@
The new function $w$ on $\mathbb{Z}^{2}$ introduced in (46) will facilitate a partial integration of the system. Eliminating $v$ the equations (46) imply,
@@
Q2 $2\alpha\beta(\alpha-\beta)$
@@
H2 $\alpha-\beta$
H1 $\alpha-\beta$
@@
System (49) indicates the function $w$ has a decomposition in potentials $U$ and $\mu$,
$w=U\frac{\mu\hat{\tilde{\mu}}}{\tilde{\mu}\hat{\mu}}.$ (51)
@@
There is an interpretation $U=\partial_{x}u$ in terms of dependence of $u$ on an independent variable $x$, except for H1 where $U=1$ and H3δ=0 where $U=u$ without loss of generality. Second, $\mu$ is a gauge factor defined by,
@@
Suppose $u$ satisfies a quad-equation in Table 1. If $\tau$ is the associated tau function, then
$U=\frac{\tilde{\sigma}\hat{\sigma}}{\sigma\hat{\tilde{\sigma}}}f_{\alpha,\beta}(u,\tilde{u},\hat{u}),\quad\sigma=\tau\mu,$ (54)
satisfies (52), where $\mu$ is a gauge factor defined by (53). Conversely, if $U$ satisfies (52), then (54) can be used to determine $\tau$.
@@
The tau function (2) of each equation in Table 1, with the exception of Q4 and A2, is a function on $\mathbb{Z}^{d}$ for $d>2$ satisfying,
$\left[\begin{array}[]{llll}0&a&b&c\\ \bar{a}&0&\bar{c}&\bar{b}\\ \tilde{b}&\tilde{c}&0&\tilde{a}\\ \hat{c}&\hat{b}&\hat{a}&0\end{array}\right]\left[\begin{array}[]{l}\tau\hat{\tilde{\bar{\tau}}}\\ \bar{\tau}\hat{\tilde{\tau}}\\ \tilde{\tau}\hat{\bar{\tau}}\\ \hat{\tau}\tilde{\bar{\tau}}\end{array}\right]=\left[\begin{array}[]{l}0\\ 0\\ 0\\ 0\end{array}\right],\quad$ (55)
@@
$a=\frac{\mu\hat{\tilde{\mu}}}{\tilde{\mu}{\hat{\mu}}},\quad b=\frac{\mu\hat{\bar{\mu}}}{\hat{\mu}{\bar{\mu}}},\quad c=\frac{\mu\tilde{\bar{\mu}}}{\bar{\mu}{\tilde{\mu}}},$ (56)
and $\mu$ is a gauge factor that satisfies (53).
Proof.
The necessity can be checked by glancing at Table 1, where the polynomials defining the mentioned equations satisfy identities,
@@
and so on. Combined with (46) they convert directly to the bilinear identities for $\tau$. The sufficiency is checked by substation into the equation (24). ∎
Recall that for construction of solutions, only a particular tau function is needed. Because it avoids square roots, the solution
$\mu=\prod_{i<j}(-1)^{{\left\lfloor{n_{i}/2}\right\rfloor}n_{j}}d(\alpha_{i},\alpha_{j})^{{\left\lfloor{n_{i}n_{j}/2}\right\rfloor}},$ (58)
of (53) might be preferred, but the higher order original condition (24) was autonomous. If
$\mu=\prod_{i<j}(-1)^{{\left\lfloor{n_{i}/2}\right\rfloor}n_{j}}\alpha_{ij}^{n_{i}n_{j}},\quad\alpha_{ij}=\sqrt{d(\alpha_{i},\alpha_{j})},$ (59)
and $(\alpha,\beta,\gamma)$ correspond to $(\alpha_{i},\alpha_{j},\alpha_{k})$ where $i<j<k$, then
$a^{2}=d(\alpha,\beta),\quad b^{2}=d(\alpha,\gamma),\quad c^{2}=d(\beta,\gamma),$ (60)
in (56). This choice of $\mu$ therefore confines the non-autonomicity in (55) to flipping the signs of terms. I have not noticed any choice that makes the system completely autonomous.
@@ 4.2 Example
For H1, $d(\alpha,\beta)=\alpha-\beta$ and $h_{\alpha}(u,\tilde{u})=1$, so that $U=1$ is a solution for (52), and the equation (54) for $\sigma$ is
$\frac{\sigma\hat{\tilde{\sigma}}}{\tilde{\sigma}\hat{\sigma}}=\tilde{u}-\hat{u},$ (61)
leading to the known relation between Hirota KdV and H1 [17] in this case. For solution (5) and tau-function (7), the resulting equation for $\sigma$,
$\sigma\hat{\tilde{\sigma}}=(p-q)\tilde{\sigma}\hat{\sigma},$ (62)
has the particular solution $\sigma=(p-q)^{nm}$.
@@ 5.2 Tau function
identically as polynomials in $y$ and $z$, where
@@
$\tau=(-1)^{{\left\lfloor{m/2}\right\rfloor}n}\mathop{\mathrm{\Theta}}\nolimits(\xi)\left(\frac{\mathop{\mathrm{H}}\nolimits(\alpha-\beta)}{\mathop{\mathrm{H}}\nolimits(\alpha+\beta)}\right)^{\left\lfloor{nm/2}\right\rfloor},$ (76)
@@ 5.3 Multidimensional extension
$u=\frac{\mathop{\mathrm{H}}\nolimits(\xi)}{\mathop{\mathrm{\Theta}}\nolimits(\xi)},\quad\tau=\mathop{\mathrm{\Theta}}\nolimits(\xi)\prod_{1\leq i<j\leq d}(-1)^{{\left\lfloor{n_{j}/2}\right\rfloor}n_{i}}\left(\frac{\mathop{\mathrm{H}}\nolimits(\alpha_{i}-\alpha_{j})}{\mathop{\mathrm{H}}\nolimits(\alpha_{i}+\alpha_{j})}\right)^{\left\lfloor{n_{i}n_{j}/2}\right\rfloor},$ (77)
@@ 5.5 $Q4^{*}$ solution
They correspond to the four possible values for $c$ determined by (87) with in the particular case $r=k\mathop{\mathrm{cn}}\nolimits(\gamma)/\mathop{\mathrm{dn}}\nolimits(\gamma)$ corresponding to the original relations (66).

### v2

@@ 1.1 Example
$u=np+mq,\quad p^{2}=\alpha,\quad q^{2}=\beta.$ (5)
@@ 3 Equations for tau
which is equivalent to the form in [5]. Bilinear relations are expected [6] but only after introducing a non-autonomous scaling factor (gauge factor). The most direct statement of the more refined higher dimensional characterisation uses an intermediary Schwarzian variable [7].
@@
Like $\tau$, the quantity $\bar{\tau}/{\underaccent{\bar}{\tau}}$ is also not invariant under fractional-linear point transformations of $u$. An invariant Schwarzian variable [9, 7] is
@@
where $w$ is any given singular solution of the quad equation. $\varphi$ in (30) is characterised by equations that are invariant under fractional linear transformations, and satisfies the lattice Schwarzian KP equation [10, 11]. If $w$ is constant, then $\varphi$ satisfies the lattice Schwarzian KdV equation, and all equations in Table 1 have a singular solution that is constant throughout $\mathbb{Z}^{d}$, with the exception of $A2$ and $Q4$. The canonical forms in Table 1 are such that the constant singular solution is $w=\infty$ in all cases, so substituting this into (30) offers an explanation for the isolation of $\tau$ leading to Proposition 3.2.
@@ 4 Reduced equations for tau
in terms of the Hirota variable,
@@
The new function $w$ introduced in (46) will facilitate a partial integration of the system. Eliminating $v$ the equations (46) imply,
@@
Q2 $4\alpha\beta(\alpha-\beta)$
@@
H2 $\beta-\alpha$
H1 $\beta-\alpha$
@@
System (49) and the dimensional skew-symmetry of (46) indicates the function $w$ has a decomposition in potentials $U$ and $\mu$,
$w=U\frac{{\underaccent{\tilde}{\mu}}\hat{\mu}}{\mu\hat{{\underaccent{\tilde}{\mu}}}}.$ (51)
@@
There is an interpretation $U=\partial_{x}u$ in terms of dependence of $u$ on an independent variable $x$, except for H1 and H3δ=0. Second, $\mu$ is a gauge factor defined by,
@@
Suppose $u$ satisfies a quad-equation in Table 1. If $\tau$ is the associated tau function and $\mu$ satisfies (53), then
$U=\frac{\tilde{\sigma}\hat{\sigma}}{\sigma\hat{\tilde{\sigma}}}\frac{f_{\alpha,\beta}(u,\tilde{u},\hat{u})}{d(\alpha,\beta)},\quad\sigma=\tau/\mu,$ (54)
satisfies (52). Conversely, if $U$ satisfies (52), then $\tau$ is determined by (54).
@@
The tau function (2) of each equation in Table 1, with the exception of Q4 and A2, is a function on $\mathbb{Z}^{d}$ for $d>2$ of the form $\tau=\sigma\nu$, where $\sigma$ on $\mathbb{Z}^{d}$ satisfies the following reduction of the Miwa equation,
$\left[\begin{array}[]{cccc}0&t(\alpha)-t(\beta)&t(\beta)-t(\gamma)&t(\gamma)-t(\alpha)\\ t(\alpha)-t(\beta)&0&-1&1\\ t(\beta)-t(\gamma)&1&0&-1\\ t(\gamma)-t(\alpha)&-1&1&0\end{array}\right]\left[\begin{array}[]{l}\sigma\hat{\tilde{\bar{\sigma}}}\\ \bar{\sigma}\hat{\tilde{\sigma}}\\ \tilde{\sigma}\hat{\bar{\sigma}}\\ \hat{\sigma}\tilde{\bar{\sigma}}\end{array}\right]=\left[\begin{array}[]{l}0\\ 0\\ 0\\ 0\end{array}\right],$ (55)
and $\nu$ on $\mathbb{Z}^{d}$ is a solution of the multiplicatively linear system
$\frac{\hat{\tilde{\nu}}{\underaccent{\tilde}{\nu}}}{\tilde{\nu}{\underaccent{\tilde}{\wh{\nu}}}}={t(\beta)-t(\alpha)},\qquad\frac{\hat{\tilde{\nu}}{\underaccent{\hat}{\nu}}}{\hat{\nu}{\underaccent{\hat}{\wt{\nu}}}}={t(\alpha)-t(\beta)},$ (56)
with $t$ as defined in (25).
Proof.
The corresponding condition on $\tau$ is,
$\left[\begin{array}[]{cccc}0&\tilde{a}&\hat{b}&\bar{c}\\ \tilde{a}&0&c&\hat{\bar{b}}\\ \hat{b}&\tilde{\bar{c}}&0&a\\ \bar{c}&b&\hat{\tilde{a}}&0\end{array}\right]\left[\begin{array}[]{l}\tau\hat{\tilde{\bar{\tau}}}\\ \bar{\tau}\hat{\tilde{\tau}}\\ \tilde{\tau}\hat{\bar{\tau}}\\ \hat{\tau}\tilde{\bar{\tau}}\end{array}\right]=\left[\begin{array}[]{l}0\\ 0\\ 0\\ 0\end{array}\right],$ (57)
@@
$a=\frac{\nu\hat{\tilde{\nu}}}{\tilde{\nu}{\hat{\nu}}},\quad b=\frac{\nu\hat{\bar{\nu}}}{\hat{\nu}{\bar{\nu}}},\quad c=\frac{\nu\tilde{\bar{\nu}}}{\bar{\nu}{\tilde{\nu}}}.$ (58)
The necessity of (57) can be checked by glancing at Table 1, where the H1 polynomial satisfies identities,
@@
and so on. Combined with (46) they convert directly to the bilinear identities for $\tau$, so, according to Proposition 3.2, must hold for the common tau function of all mentioned equations. That relations (57) imply (24) can be checked by calculation using the defining properties of $\nu$, in particular that $\hat{\tilde{a}}=-a$, $\tilde{a}=-\hat{a}$, $a\tilde{a}=t(\beta)-t(\alpha)$, $\bar{a}=a$, and so on. ∎
In particular, given any tau function for the mentioned equations, there is a unique gauge factor, $\nu$ satisfying (56), such that $\sigma=\tau/\nu$ satisfies (55). Furthermore, linear contraction on (57) acts trivially on $\nu$, so descends to an action on $\sigma$ as follows (constants a and b are absourbed into $\nu$).
Proposition 4.3.
If $\sigma$ satisfies (55) on $\mathbb{Z}^{d+1}$, then
$\sigma^{\prime}=\left.\frac{\tilde{\nu}\tilde{\sigma}+{\underaccent{\tilde}{\nu}}{\underaccent{\tilde}{\sigma}}}{\nu}\right|_{\mathbb{Z}^{d}}$ (60)
satisfies (55) on $\mathbb{Z}^{d}$, where $\nu$ on $\mathbb{Z}^{d+1}$ is any solution of (56).
The linear contraction on $\tau$ has the advantage of being cleanly separated from the symmetries (3) and (4) that alter $\nu$, but in this form on $\sigma$ it can be recognised as a Bäcklund transformation for the Miwa equation that is compatible with the reduction.
@@ 4.2 Example
A solution of the Miwa equation reduction (55) is
$\sigma=\prod_{i<j}(p_{i}+p_{j})^{-n_{i}n_{j}},\qquad t(\alpha_{i})=p_{i}^{2},\quad i,j\in\{1,\ldots,d\}.$ (61)
This solution is related to the ongoing tau function example (8),
$\tau=\prod_{i<j}(-1)^{{\left\lfloor{n_{i}/2}\right\rfloor}n_{j}}\left(\frac{p_{i}-p_{j}}{p_{i}+p_{j}}\right)^{{\left\lfloor{n_{i}n_{j}/2}\right\rfloor}},$ (62)
because $\nu=\tau/\sigma$ satisfies system (56) with $t(\alpha_{i})=p_{i}^{2}$.
@@ 5.2 Tau function
identically as a polynomial in $y$ and $z$, where
@@
$\tau=(-1)^{{\left\lfloor{n/2}\right\rfloor}m}\mathop{\mathrm{\Theta}}\nolimits(\xi)\left(\frac{\mathop{\mathrm{H}}\nolimits(\alpha-\beta)}{\mathop{\mathrm{H}}\nolimits(\alpha+\beta)}\right)^{\left\lfloor{nm/2}\right\rfloor},$ (76)
@@ 5.3 Multidimensional extension
$u=\frac{\mathop{\mathrm{H}}\nolimits(\xi)}{\mathop{\mathrm{\Theta}}\nolimits(\xi)},\quad\tau=\mathop{\mathrm{\Theta}}\nolimits(\xi)\prod_{1\leq i<j\leq d}(-1)^{{\left\lfloor{n_{i}/2}\right\rfloor}n_{j}}\left(\frac{\mathop{\mathrm{H}}\nolimits(\alpha_{i}-\alpha_{j})}{\mathop{\mathrm{H}}\nolimits(\alpha_{i}+\alpha_{j})}\right)^{\left\lfloor{n_{i}n_{j}/2}\right\rfloor},$ (77)
@@ 5.5 $Q4^{*}$ solution
They correspond to the four possible values for $c$ determined by (87) in the particular case $r=k\mathop{\mathrm{cn}}\nolimits(\gamma)/\mathop{\mathrm{dn}}\nolimits(\gamma)$ corresponding to the original relations (66).